Divide 1161 underpower 8 with 31 underpower 8. Express the obtained result in decimal number.
1
Expert's answer
2019-03-05T12:28:54-0500

There is one easy way to solve this.

Convert "1161_8" to the decimal system by adding each digit from left to right multiplied by 8 in power of {position (counted from the rightmost digit) minus 1}:


"1161_8=1\cdot 8^{4-1}+1\cdot 8^{3-1}+6\cdot 8^{2-1}+1\cdot 8^{1-1}=625,"

"31_8=3\cdot 8^1+1\cdot 8^0=25,"

divide them in the decimal form (625/25), get 25, which as we saw above is "31_8".

There is another way to solve this applying column division, but it is more complex, takes more time and is not easy to understand.